OpenStudy (anonymous):

-2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation?

OpenStudy (whpalmer4):

A natural next step would be to apply the distributive property to remove the parentheses.

OpenStudy (whpalmer4):

You could also add 7 to both sides.

OpenStudy (anonymous):

help An equation is shown below: -2(4x - 1) - 7 = 5 Which statement shows a correct next step in solving the equation? The equation can become -2(4x - 1) = -2 by applying the distributive property. The equation can become -2(4x - 1) = 12 by applying the addition property of equality. The equation can become -2(4x - 1) = 12 by applying the commutative property of addition The equation can become -2(4x - 1) = -2 by applying the subtraction property of equality.

OpenStudy (whpalmer4):

Okay, I'll give you another hint: we added 7 to both sides.

OpenStudy (whpalmer4):

Look at your book and see which of those properties has a definition that would support adding 7 to both sides of the equation.

OpenStudy (whpalmer4):

You need to know all of those properties to make any headway in algebra, so it is time well spent learning them so they become second nature to you.